if two opposite sides of a square are each increased by 12 inches, and the other two sides are increased 4 inches, a rectangle is formed having a perimeter double that of the square. how long are the sides of the square?
Answer by ankor@dixie-net.com(22740)

if two opposite sides of a square are each increased by 12 inches,
and the other two sides are increased 4 inches,
a rectangle is formed having a perimeter double that of the square.
how long are the sides of the square?
:
Let x = the side of the square (4x = the perimeter)
then
(x+12) = the length of newly formed rectangle
and
(x+4) = the width of the rectangle
:
the equation for the statement:
" a rectangle is formed having a perimeter double that of the square."
2(x+12) + 2(x+4) = 2(4x)
simplify, divide by 2
(x+12) + (x+4) = 4x
2x + 16 = 4x
16 = 4x - 2x
2x = 16
x = 16/2
x = 8 in is the side of the square
:
;
You can confirm this by finding the perimeters of each, see if the rectangle
is twice the square